The following information pertains to the Moline Facility for the month of May (all materials are added at the beginning of t 0% completion
Answer #1

1) Total Cost to be accounted for = 28000+110400 = 138400

Equivalent unit = 110000*0%+120000+140000*100%) = 260000

Cost per equivalent unit = 110400/260000 = 0.42

2) Total Cost to be accounted for = 32000+100400 = 132400

Equivalent unit = 80000+280000 = 360000

Cost per equivalent unit = 132400/360000 = 0.37
Answer #2

1) units                    materials

   110,000               28,000

  260,000                110,400

  230,000

  140,000


step 1) 110,000+260,000= 370,000


230,000- 140,000= 90,000

physical units

110,000*100%= 110,000          90,000*100%= 90,000

140,000*100%= 140,000

total= 110,000+ 140,000+ 90,000= 340,000


materials 

0

90,000

140,000

230,000


step 2

cost to be acct for answer: 110,400

cost per equivalent( 110,400/230,000)= 0.48